OpenStudy (lukecrayonz):

The path of a softball is given by the equation: y=-0.08x^2+x+4 The coordinates x and y are measured in ft, with x=0 corresponding to the position where the ball was thrown. (a) use a graphing utility to graph the trajectory of the softball (b) Move the cursor along the path to approximate the highest point and the range of the trajectory

OpenStudy (lukecrayonz):

http://screensnapr.com/v/WEXz8m.png

OpenStudy (lukecrayonz):

Seem good?

OpenStudy (lukecrayonz):

Oh hey blossom, my girlfriend just bought me one direction tickets haha

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D-0.08x%5E2%2Bx%2B4+ Highest point is y=57/8 at x=25/4. The range looks like its around 15.6 ft.